Circular No. 7241

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999dn IN NGC 7714 Y. L. Qiu, Q. Y. Qiao, and J. Y. Hu, Beijing Astronomical Observatory

(BAO), report their discovery, in the course of the BAO Supernova Survey, of an apparent supernova (mag ∼ 16.0) in NGC 7714. SN 1999dn was discovered on two unfiltered CCD images taken on Aug. 19.76 and 19.82 UT with the BAO 0.6-m telescope, and it was confirmed on Aug. 20.65 with the BAO 0.60-m Schmidt telescope by X. Zhou and Z. Zheng. The new star is located at α = 23^{h}36^{m}14^{s}.70, δ = +2^{o}09′08′′.8, which is 9′′.9 east and 9′′.4 south of the center of NGC 7714. This is a bright foreground star near NGC 7714. CCD frames taken on Aug. 10 (limiting mag ∼ 18.0) show no object at the position of the supernova.

SUPERNOVA 1999dm IN MCG +05-36-022 Corrigenda. On IAUC 7237, the position for the nucleus of the host

galaxy (MCG +05-36-022 = Abell 2065-164) of SN 1999dm was provided, rather than the supernova’s position. Lee et al. provide the supernova’s position as α = 15^{h}22^{m}28^{s}.90, δ = +27^{o}42′58′′.0 (equinox 2000.0). The reference on line 8 was published in 1988, not 1998. On line 11, for BAO reflector read BOA reflector

M. M. M. Santangelo and S. Donati, Monte Agliale, Italy, report that their CCD patrol frames of the galaxy cluster Abell 2065 obtained on Mar. 24.09 and May 17.96 UT with the 0.51-m reflector (limiting mag R ∼ 20.0) show nothing new near the host galaxy.

COMET C/1999 N2 (LYNN)

Ephemeris from orbital elements on MPC 35208:

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Aug. 10 12^{h}22.^{m}90 +20^{o}27.^{′}9 1.178 0.838 44.^{o}1 57.^{o}4  7.6
15 12 32.38 +23 07.2 1.305 0.882 42.4 50.8  8.0
20 12 40.06 +25 17.1 1.425 0.933 40.8 45.1  8.5
25 12 46.58 +27 05.9 1.535 0.989 39.4 40.4  8.9
30 12 52.37 +28 39.5 1.637 1.050 38.2 36.5  9.3
Sept. 4 12 57.73 +30 02.2 1.729 1.114 37.5 33.5  9.7
9 13 02.84 +31 17.2 1.812 1.179 37.2 31.1 10.0
14 13 07.85 +32 27.2 1.887 1.247 37.4 29.3 10.3
19 13 12.82 +33 34.1 1.952 1.315 38.0 28.1 10.6
1999 August 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green